If the MPC is 0.95 and the MPM is 0.05, the open economy multiplier is 10.

Select one:

True

False
Answer #1

True

Reason

Open economy multiplier = 1/ MPS + MPM

We know that MPS = 1-MPC

So MPS = 1- 0.95 = 0.05

So Multiplier = 1/0.05+0.05 =1/0.10 = 10